OpenStudy (anonymous):

Find 3 consecutive odd numbers where the product of the smaller two numbers is 46 less than the square of the largest number.@jiraya @alli14344

OpenStudy (anonymous):

Let x be the smallest of the three numbers. the numbers would be x, x+2, x+4 (understand why?) so the product of the smallest two would be: x(x+2) 46 less than the square of the largest would be: (x+4)^2 - 46 set these equal and solve. x(x+2) = (x+4)^2 - 46

OpenStudy (anonymous):

5, 7, 9 Let X be the first odd number, so X+2 will be the next odd number, and X+4 will be the final odd number. X * (X+2) = (X+4)2 - 46 X2+2X = X2+8X+16-46 X2+2X = X2+8X-30 Rearrange to get like terms to one side, 6X = 30 X =5 So the next two odd numbers would be 7 & 9 We can check the results: 5*7 =? 92 -46 35 =? 81-46 35 = 35 So the results are correct.